# Conceptualizing Trustworthiness and Trust in Communications

### v1

@@ IV-A Related Work
Similarly, Veith et al. published a survey paper analysing potential trust anchor technologies for 6G systems [11]. Besides technologies, they also analysed the relationship between technology and trust concepts as developed by humanities. Thereby they also refer to the concepts of ability, integrity, and benevolence. They also identified the question of quantifying and measuring trustworthiness as an important yet unanswered one. As in the NGMN report, no clear distinction between trustworthiness and trust is made.
@@ V A model of trust and trustworthiness in the relationship between humans and technology
Given the concepts of trust and trustworthiness introduced in the previous sections, we now develop a unifying model which integrates both concepts. We start with a basic model, which we will extend in the next section.
@@ V-A Basic Model integrating Trust and Trustworthiness
Although TCs can be measured in principle, especially “absolute” measurements of TCs must be put into context and are only a prerequisite for making a trust decision. Additionally, emotions play a role in human decisions, which in technical terms are related to a combination of general long-term experience and acute experience/situation context.
So, the selection and weighting of the objective measurable TCs in decision-making is context and experience-dependent. And, based on the experience after decision-making and action-taking, information is fed back into the weighting for the next move. This feedback loop needs to be modelled. It depends on key factors that play a role in deciding on trust which are:
@@
    Own experience & learnings (subjective)
@@
In an ideal situation, humans should take objective measurements of trustworthiness characteristics (TCs), the current situation (our state), as well as our learnings into account, to make an informed decision on what to trust and to which extent. Eventually, our observed experience based on the decision will be added to our memory of learnings. So, the available TCs should be taken into account, filtered and weighted depending on the state, as well as depending on the memory of learnings. And the learnings again depend on the observations and experiences from previous similar situations. Fig. 2 shows this basic feedback loop as a block diagram.
@@
Note that, in reality, the trust decisions of a human being might only be a little influenced by objective TC measurements. Indeed, the subjective weighting of a human might just end up in deciding on the green car instead of the blue one – not only because they like green more than blue but because the colour might emotionally lead to trust and trust-related decisions. Remember that according to Mayer et al., trust-based decisions are related to dealing with risks for the trusting party. Obviously, the colour of the car will only have negligible influence on the safety of the car and therefore the foundation of such a trust decision might not contribute to the intended risk reduction – even if it might contribute to a reduction of the perceived risk. Therefore, it is of utmost importance that systems which can harm humans are designed and operated following the “trustworthiness by design/default” paradigm. Then humans can indeed make subjective trust decisions and will still not face unacceptable risks thanks to the safeguards achieved through the objective TCs. Thus, certain TCs must not be up for human weighting in a proper design. This opens up a societal and regulatory dimension.
@@ V-B Extended model: Mediators for dealing with complexity
Making trust decisions based on TCs poses the challenge of measuring them and putting them in relation to own expectations. People might not be capable or willing to do this on their own. Instead, such measurement can be delegated to third parties in the form of reviews, certification, or audits. We add these third parties as mediators in our model (Fig. 3).
@@
The concept of mediators induces a recursive challenge to trustworthiness: The trustworthiness of the mediator, e.g. a certification body, influences the trustworthiness of the mediator result (e.g. certification). On the one hand, we can apply well-known organizational and regulatory means to lower the risk of misbehaving mediators shifting the problem from a purely technical one to a more social sciences-related one. On the other hand, the design of the technical system should support as much as possible the trustworthiness assessment by every human. Although we stated that most people will not have the ability to do this assessment, some will and can act as an additional safeguard concerning misbehaving mediators.
@@
Related to the concept of mediators is the concept of reputation. Here basically other humans act as mediators giving their judgement concerning trustworthiness and their past trust-related experiences. Imagine e.g. a scenario from the domain of autonomous vehicles used in public transportation. Past users can make a judgement regarding their experience of using a given autonomous vehicle. This kind of reputation can be incorporated by future users regarding their trust decisions.
VI Auditability and Reputation as Two Meta-Requirements for Trustworthy Systems
@@
Regarding the design of a technical system, we can derive two fundamental requirements (from which subsequent requirements can be derived) from the trust model of Fig. 3. We call them meta-requirements as they are on a completely different, much more abstract, level than any of the functional or non-functional requirements we discussed so far. The first requirement is the support of auditability and the second is the support/integration of reputation systems.
Support of auditability has many technical implications. First of all, it would require that the design and specifications of the system be accessible to the mediator. Looking at the mobile networks’ ecosystem, the current attempt by the O–RAN ALLIANCE to create open specifications for the radio access network of future mobile networks is a first step in the right direction. Second, it would require that some meaningful judgements (e.g. objective measurements) regarding relevant TCs and therefore the overall trustworthiness are possible. This would require, e.g. (formal) specifications which would allow formal reasoning to support the required level of verification and assurance. Looking at the current style of mobile network specifications e.g. produced by the 3rd Generation Partnership Project (3GPP), we can identify a huge gap, since the current specifications are mainly written in natural language with some kind of semi-formal formulations at best. Besides verifiable specifications, one would also need means to assess the implementation at any given time during the operation of the system. To this end, means such as remote attestation and reproducible builds are necessary for ensuring that an executed component matches the audited source code. To support the reasoning even more, means for strict separations are necessary allowing to break down the overall assessment into smaller and independent pieces.
Support for reputation requires that the reputation system itself is trustworthy. This in turn requires some means of accountability to avoid fake reputation values. Introducing accountability could lead to reduced privacy. Imagine again the autonomous vehicle scenario described above. If the reputation system had been designed in a way that every judgement of a user is linkable to the identity of that user, then such a reputation system could reveal usage patterns and movement profiles of the users. Therefore, incorporating reputation systems into the overall technical system to support the judgements regarding trustworthiness also requires a system design approach that addresses the reputation challenge in its entirety.
These non-exhaustive considerations illustrate that the design of a trustworthy system needs to be approached holistically with support for auditability by mediators as well as for reputation integrated from the very beginning.
@@ VII Conclusion
In this paper, we presented a model which strives to clarify the relationship between “trustworthiness” and “trust” in the interaction of humans with technology. It is important to understand that trustworthiness and trust are not synonyms but have fundamentally different meanings. Trustworthiness refers to objective characteristics of a given technical system whereas trust relates to subjective behaviour or beliefs of humans. Ideally, trustworthiness influences the trust decisions in a way which reduces the risks for human beings. This concept could possibly be extended to machine-to-machine interactions.
@@
Nevertheless, a trustworthy system design is of utmost importance – especially regarding the upcoming 6G communications system that will connect an unseen number of physical objects we deal with every day. This challenge must be approached in a holistic manner supporting auditability and reputation systems.
The paradigm of “trustworthiness by design/default” requires not just security technologies to be incorporated. If truly embraced, it will change system design very fundamentally.
Acknowledgement

### v2

@@ IV-A Related Work
Similarly, Veith et al. published a survey paper analyzing potential trust anchor technologies for 6G systems [16]. Besides technologies, they also analyzed the relationship between technology and trust concepts as developed by humanities. Thereby they also refer to the concepts of ability, integrity, and benevolence. They also identified the question of quantifying and measuring trustworthiness as an important yet unanswered one. As in the NGMN report, no clear distinction between trustworthiness and trust is made.
Bauer [3] analyzes the relation between trust and trustworthiness from a social science perspective. He agrees with our view of the nature of trust being a subjective one. Our approach is different insofar as we regard trustworthiness an objective characteristic of a technical system.
@@ V A Model of Trust and Trustworthiness in the relationship between Humans and Technology
Based on the concepts of trust and trustworthiness introduced in section II and measurable TCs in section III, we now develop a unifying model which integrates both concepts. We start with a basic model, which we will extend in the next section.
@@ V-A Basic Model integrating Trust and Trustworthiness
Although TCs can be measured in principle, above all, “absolute” measurements of TCs must be put into context and are only a prerequisite for making a trust decision. Additionally, emotions play a role in human decisions, which, in technical terms, are related to a combination of general long-term and acute experiences (i.e., situation context).
Hence, the selection and weighting of the objective measurable TCs in the decision-making process depends on context and experience. In addition, the decision-making is based on the accumulated experiences collected from previous decisions and actions as well as the current experience (emotional state), as shown by empirical studies (e.g., [1]). This forms a feedback loop of experiences that needs to be modelled as well. It depends on a couple of key factors that play a role in making a trust decision. These key factors are:
@@
    Personal experience & learnings (subjective)
@@
In an ideal situation, humans should take objective measurements of trustworthiness characteristics (TCs), the current situation (one’s own state), as well as their learnings into account to make an informed decision on what to trust and to which extent. Eventually, the observed experience based on the decision will be added to the memory of learnings. Thus, the available TCs should be taken into account, filtered, and weighted depending both on the state and on the memory of learnings. In turn, the learnings depend on the observations and experiences from previous similar situations. Fig. 2 shows this basic feedback loop as a block diagram.
Note that, in reality, the trust decisions of a human being might only slightly be influenced by objective TC measurements. Indeed, people might just end up deciding on the green car instead of the blue one, for example, as a result of their subjective weighting – not only because they like green more than blue but because the color might emotionally lead to trust and trust-related decisions. Remember that according to Mayer et al., trust-based decisions are related to dealing with risks for the trusting party. Obviously, the color of the car will only have negligible influence on the safety of the car, and, therefore, the foundation of such a trust decision might not contribute to the intended risk reduction – even if it might contribute to a reduction of the perceived risk. Therefore, it is of utmost importance that systems which can harm humans are designed and operated following the “trustworthiness by design/default” paradigm. In this way, humans can indeed make subjective trust decisions and still will not face unacceptable risks, thanks to the safeguards achieved through the objective TCs. Thus, in a proper design, certain TCs must not be left to human weighting. This opens up a societal and regulatory dimension.
@@ V-B Extended model: Mediators for dealing with complexity
Making trust decisions based on TCs poses the challenge of measuring them and putting them in relation to one’s own expectations. People might not be capable or willing to do this on their own. Instead, such measurements can be delegated to third parties in the form of reviews, certification, or audits. We add these third parties as mediators in our model (Fig. 3).
@@
The concept of mediators induces a recursive challenge to trustworthiness: The (organizational) trustworthiness of the mediator itself, e.g., of a certification body, influences the (rational) trustworthiness of the result of mediator’s concrete action (e.g., certification). On the one hand, we can apply well-known organizational and regulatory means to lower the risk of potentially misbehaving mediators, thus shifting the problem from a purely technical one to a more social sciences-related one. On the other hand, the design of the technical system should support the trustworthiness assessment by every human in the best possible way. Although we stated that most people will not have the ability to do this assessment, some will and can act as an additional safeguard with regard to misbehaving mediators.
@@
Related to the concept of mediators is the concept of reputation. Here, basically, other humans act as mediators, giving their judgement concerning trustworthiness and their past trust-related experiences. Imagine, for example, a scenario from the domain of autonomous vehicles used in public transportation. Past users can make a judgement based on their experience of using a given autonomous vehicle. The ensuing reputation can then be taken into account by future users when facing trust decisions.
@@
To support the role of mediators as well as the concept of reputation, trustworthy technical systems must fulfill two fundamental requirements (from which subsequent requirements can be derived). We call them meta-requirements, as they are on a completely different, much more abstract, level than any of the functional or non-functional requirements we discussed so far. The first requirement is the support of auditability, as needed by the mediators, and the second is the support/integration of reputation systems.
VI Thoughts on an Empirical Evaluation of the Proposed Model
We derived our trust model from the definition of trustworthiness and the existing ABI model. In the following, we want to justify it by finding empirical evidence that a change in the metrics of the objective TCs indeed influences the trust decisions of humans. A more specific example would be to show that an increase in the measured values of security- and safety-related TCs indeed increases the likelihood that the trust decisions of humans are geared towards systems with higher objective TC values. We selected security- and safety-related TCs as examples because we assume that, if all other things are equal, humans prefer systems which offer higher security and safety. On the other hand, our model implies that not all humans will automatically decide for the system which provides higher security and safety – because the subjective weighting factors influence the final outcome of the trust decision.
For example, the study by Wiedemann et al. [18] supports this approach. It analyzes the influence of the specific absorption rate (SAR) with regard to customer purchase decisions for mobile phones. Here, we can understand the SAR value as one objective TC, and the purchase decision as the trust decision made by humans. The study results showed that customers perceive higher SAR values as a greater risk, leading to purchasing decisions towards phones with lower SAR values. This can be interpreted as an objective TC which influences the trust decisions of the customers. On the other hand, the study showed that a lower SAR value of a given phone does not automatically lead to a purchase decision for that phone – this could be seen as an indicator that subjective weighting factors eventually influence the trust decision.
Other examples are the various battery and overheating problems of mobile phones, as with the Motorola Razr (2005), HTC One M9 (2015) or Samsung Galaxy Note 7 (2016). In all cases, these problems led to a decrease in sales, which again can be seen as some evidence for the claim that a change (i.e., decrease) in an objective (safety-related) TC influences trust-based decisions (i.e., sales). These examples also show the importance of the mediators since, in fact, the objective TC had not changed – only the knowledge about it, thanks to reports in the media.
On the other hand, there are cases in which low metrics of certain objective TCs become known, and such knowledge does not lead to changing trust decisions, nor, specifically, to decreases in sales either. We could attribute this behavior of the consumers to the subjective weighting factors of our model. One example is the iPhone 6 “Bendgate” (2014), which came to be known to be susceptible to chassis bending. Nevertheless, the iPhone 6 achieved sales records. A second example is the HTC One M8 overheating issue (2014), which did not lead to notable decreases in sales.
A third example can be taken from the vehicular domain: Adaptive cruise control (ACC) systems enable a vehicle to automatically keep a minimum distance to vehicles in front. From a technical, i.e., safety, related point of view, it would be sufficient to have a fixed, predefined minimum distance (and which is potentially speed dependent). Nevertheless, many car manufacturers allow users to configure a larger distance than necessary. In this example, the trust-related decision is to enable/use the ACC system. Given a certain type of car, the objective TCs related to the ACC system are basically equal for all users. Therefore, the option to configure the minimum distance accounts for the subjective weighting factors of individual drivers.
@@ VII Conclusion
In this paper, we presented a model which strives to clarify the relationship between “trustworthiness” and “trust” in the interaction of humans with technology. It is important to understand that trustworthiness and trust are not synonymous but have fundamentally different meanings. Trustworthiness refers to objective characteristics of a given technical system, whereas trust relates to subjective behavior or beliefs of humans. Ideally, trustworthiness influences the trust decisions in a way which reduces the risks for human beings. This concept could possibly be extended to machine-to-machine interactions.
@@
Nevertheless, a trustworthy system design is of utmost importance – especially regarding the upcoming 6G communications system that will connect an unheard-of number of physical objects we deal with every day. This challenge must be approached in a holistic manner supporting auditability and reputation systems.
The paradigm of “trustworthiness by design/default” does not just call for security technologies to be incorporated. If truly embraced, it will change system design very fundamentally.
Acknowledgements